OpenStudy (anonymous):

Simplify the expression. sin^2x-1/cos(-x)

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

is that \(\Large \dfrac{\sin^2x-1}{\cos (-x)}\)

OpenStudy (anonymous):

yes

hartnn (hartnn):

do you know \(\sin^2x + \cos^2x = 1\) ?

OpenStudy (anonymous):

thank you both